OpenStudy (anonymous):

Find the point on the curve y=x^2 where the slope the tangent is equal to the x coordinate of that point.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

y=x^2 dy/dx=2x also dy/dx=x 2x=x 2x-x=0 x=0 Hence y=0^2=0 point is (0,0)